What is the velocity and direction of the collision if a rail way train with mass of 6000kg is moving to

the right at 2.0m/s and collides with an empty freight train mass 3,000kg which is moving to the left on

the same track at 3.0m/s?


1
Expert's answer
2021-11-22T10:12:35-0500

We can find the velocity of the combination of railway train and freight train from the law of conservation of momentum:


"m_1v_1-m_2v_2=(m_1+m_2)v_c,""v_c=\dfrac{m_1v_1-m_2v_2}{m_1+m_2},""v_c=\dfrac{6000\ kg\times2.0\ \dfrac{m}{s}-3000\ kg\times3.0\ \dfrac{m}{s}}{6000\ kg+3000\ kg}=0.33\ \dfrac{m}{s}."

The sign plus means that the combination of railway train and freight train after the collision moves to the right.